# A Nonparametric Bayesian Local-Global Model for Enhanced Adverse Event Signal Detection in Spontaneous Reporting System Data

### v2

@@ 3.2 The Gamma Poisson Shrinker (GPS) model
where $\kappa_{ij}=\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})/[\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})+(1-\kappa)f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{2},\beta_{2})]$. Under an empirical Bayes approach, the point estimates $\hat{\kappa}$, $\hat{\alpha}_{1}$, $\hat{\beta}_{1}$, $\hat{\alpha}_{2}$, and $\hat{\beta}_{2}$ for the hyperparameters are obtained by maximizing their marginal likelihoods given $n_{ij}$ and $E_{ij}$. The resulting posterior of $\lambda_{ij}$ conditional on these estimated hyperparameters is used for inference and signal detection. Specifically, dumouchel2001empirical define the empirical Bayesian score $E(\log_{2}(\lambda_{ij})\mid n_{ij},E_{ij},\hat{\kappa},\hat{\alpha}_{1},\hat{\beta}_{1},\hat{\alpha}_{2},\hat{\beta}_{2})$ as a point estimate of signal strength for the drug-$j$ and AE-$i$ pair, and also suggest determining the pair to be a signal if the 5th percentile (EB05) of the above conditional posterior for $\lambda_{ij}$ exceeds a specific threshold, commonly 2. Similar to the BCPNN method, the deterministic threshold of $2$ may not necessarily ensure strong FDR control; however, a data-driven threshold controlling FDR to a pre-specified level (e.g., 0.05) may be obtained.ahmed2010false
3.3 DP Poisson method of Hu et al.hu2015signal
Several extensions and generalizations of the two-component mixture prior used in the GPS method have been suggested in the literaturehu2015signal ; zhang2018three ; tan2025flexible . Of these, we focus on one of the most flexible generalizations as proposed in Hu et al.hu2015signal , which employs a Dirichlet process mixture as the prior for $\lambda_{ij}$. Hu et al.hu2015signal consider the same Poisson likelihood $n_{ij}\sim$ Poisson($\lambda_{ij}E_{ij}$) as in GPS, but replace the two-gamma mixture prior for $\lambda_{ij}$ by a DP (infinite) mixture prior:
@@
Meanwhile, the atoms $\theta_{hj}$ are generated independently from the DP base gamma distribution: $\theta_{hj}\sim G(\beta_{j})$; $h\geq 1$. In practice, the infinite mixture (1) defining $D_{j}$ can be well-approximated ishwaran2002exact by a finite mixture $\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$ with a sufficiently large number of components $K$; this finite mixture approximation provides a computationally tractable approach to implementing this model.
Hu et al.hu2015signal consider full Bayesian inference on $\lambda_{ij}$ by first eliciting independent vague prior distributions on the hyperparameters $\alpha_{j}$ (specifically a Uniform(0.2, 10) prior) and $\beta_{j}$ (a Uniform(0, 1) prior) and then utilizing MCMC sampling from the resulting joint posterior. Subsequently, the drug-$j$ and AE-$i$ pair is identified as a signal if the 5th posterior percentile of $\lambda_{ij}$ computed using its posterior MCMC draws exceeds a certain threshold. The authors suggest using the deterministic threshold of $2$ for this purpose; a more data-driven threshold controlling FDR at a prespecified level (e.g., $0.05$) may be obtained.ahmed2010false
@@ 3.4 ZIP model based Pseudo LRT (frequentist) method
In parallel to Bayesian developments, major advances on SRS data mining for AE signal detection have also been made using frequentist methods, notably using the machinery of the likelihood ratio test huang2011likelihood ; huang2013likelihood ; huang2014likelihood . These approaches often have the advantage of built-in FDR control while ensuring high sensitivity. We consider the ZIP model for the reporting count $n_{ij}\sim\text{ZIP}(n=n_{ij}\mid\lambda_{ij}\times E_{ij},\omega_{j})$. The pseudo LRT, as proposed by Chakraborty et al.chakraborty2022use , is a recent method designed to accommodate the zero-inflated nature of SRS data while easing computational demands. Similar to the standard LRT method, this approach first tests a global alternative hypothesis $H_{aj}:\lambda_{ij}>1$ for at least one $i\in\{1,\dots,I\}$ against the null hypothesis $H_{0j}:\lambda_{ij}=1$ for all $i\in\{1,\dots,I\}$ for a given drug-$j$. If $H_{aj}$ is rejected, a step-down procedure is used to conduct individual tests on $H_{aij}:\lambda_{ij}>1$.
@@
The maximum likelihood ratio statistic for drug-$j$ is then defined as $\text{MLR}_{j}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\}\end{subarray}}\text{LR}_{ij}$. In the global test, a parametric bootstrap is used to approximate the p-value $P_{j}=\Pr(\text{MLR}_{j}\geq\text{MLR}_{j}^{\text{obs}})$ using (parametric bootstrap) draws for $\text{LR}_{ij}$ under $H_{0j}$ and hence $\text{MLR}_{j}$, where $\text{MLR}_{j}^{\text{obs}}$ is the observed test statistic. If the global test is rejected, step-down p-values chakraborty2022use ; huang2011likelihood are computed for the individual tests using the same parametric bootstrap samples for $\text{MLR}_{j}$ and using the individual observed likelihood ratio test statistic values, such that $P_{ij}=\Pr(\text{LR}_{ij}\geq\text{MLR}_{ij}^{\text{obs}})$. For multiple drugs, the same approach can be applied, defining an overall maximum likelihood ratio as $\text{MLR}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\},j\in\{1,\dots,J\}\end{subarray}}\text{LR}_{ij}$. This step-down procedure employing the same global test statistic strictly controls the false discovery rate at the same level $\alpha$ of the global test performed using the MLR huang2011likelihood ; chakraborty2022use . When multiple drugs $j$ are tested simultaneously, one considers the extended version with global maximum likelihood ratio $\max_{j\in\{1,\dots,J\}}\text{MLR}_{j}$.
@@ 4 A NOVEL LOCAL-GLOBAL DP MODEL FOR SRS DATA MINING
We begin this section by revisiting the DP Poisson model of Hu et al. hu2015signal (Section 4.1), highlighting the role of its hyperparameters in SRS signal discovery with the aid of a schematic plot (Figure 2a) and providing guidance on prior specification for these hyperparameters. We then introduce our novel local-global DP framework (Section 4.2), which accommodates between-drug associations and enables effective information sharing across drugs, thereby enhancing SRS data mining.
@@
The stick-breaking representation (2) elucidates that when $\alpha_{j}<1$, each $v_{hj}$, on average, is greater than $1/2$ with smaller values of $\alpha_{j}$ favoring $v_{hj}$ closer to one. Consequently, $\eta_{hj}$ decreases sharply in $h$, leading the mixture (1), or its finite approximation with $K$ components, to retain only a few non-empty components with sufficiently large and positive weights $\eta_{hj}$. Under a finite mixture approximation with $K_{+}<K$ non-empty components, (1) models $\lambda_{1j},\dots,\lambda_{Ij}$ using a categorical distribution with $K_{+}$ categories, placing point masses at the atoms $\theta_{1j},\dots,\theta_{K_{+}j}$ with weights $\eta_{1j},\dots,\eta_{K_{+}j}$. This fosters clustering, restricting $\lambda_{1j},\dots,\lambda_{Ij}$ to at most $K_{+}$ distinct values. When $K_{+}<I$, multiple $\lambda_{ij}$ share the same cluster $h$ admitting the common value $\theta_{hj}$, facilitating their pooling for information sharing. It follows that a smaller $\alpha$ encourages a smaller $K_{+}$, resulting in fewer clusters and thus greater information sharing among the $\lambda_{ij}$, while a larger $\alpha$ induces a larger $K_{+}$, resulting in more clusters that collectively capture more heterogeneity among the $\lambda_{i}j$ values. Importantly, when supported by the data and induced through a (moderately) large $\alpha$, values of $K_{+}>2$ are permitted in the posterior of this method. This contrasts with the GPS method of dumouchel1999bayesian (see Section 3 for a review) which strictly categorizes $\lambda_{ij}$ into two groups, intuitively, signal ($\lambda_{ij}\leq 1$) and non-signal ($\lambda_{ij}\gg 1$). The DP framework, by allowing $K_{+}>2$, enhances heterogeneity acknowledgment within both signal and non-signal groups. Figure 2a illustrates this with $I=6$ AEs and $K_{+}=3$ clusters.
@@
Because a smaller $\alpha_{j}$ promotes greater information sharing, a prior assigning sufficient mass to small positive values is recommended for stable estimation of $\lambda_{ij}$. Hu et al.hu2015signal used a Uniform(0.2, 10) prior for $\alpha_{j}$. We suggest a more aggressive weakly informative default prior, viz., Exponential(rate = $\psi_{\alpha}$) with a moderately large $\psi_{\alpha}$, e.g., $\psi_{\alpha}=3$ (placing $\approx 95\%$ mass in $(0,1)$). This prior favors small $\alpha_{j}$ while allowing larger values if needed. This choice also ensures conjugacy for efficient Gibbs sampling updates for MCMC implementation.
@@
Figure 2: Schematic diagrams visualizing the mechanisms of a “local only” DP prior with one drug (Panel A) and local-global mixture DP prior with $J=3$ drugs (Panel B). Throughout, darker shades represent larger values. Panel A: The observed relative reporting rates $n_{ij}/E_{ij}$ for $I=6$ AEs for the drug $j$ (Panel A.1) show high variability across $i$ (different shades of green) due to randomness in observed data. The atoms $\theta_{hj}$ (Panel A.2) of the DP prior enable clustering of $I=6$ AE signal strengths into $K_{+}=3$ non-empty DP clusters. This ensures information sharing: the AEs $(i=1,i=3,i=5)$ in cluster 1 and separately the AEs $(i=2,i=6)$ in cluster 2 (cluster memberships are visualized via arrows between Panels A.1 and A.2) share their observed $(n_{ij},E_{ij})$ information to inform their cluster-specific DP atoms (single atom per cluster). These DP atoms, together with the cluster memberships, produce the final $I=6$ AE signal strengths $\lambda_{ij}$ (Panel A.3). Panel B: The noisy observed data $(n_{ij},E_{ij})$ leading to noisy observed relative reporting rates $n_{ij}/E_{ij}$ (Panel B.1; different shades of green) are used to produce “local” signal strengths $\lambda_{ij}^{l}$ for $I=6$ AEs and $J=3$ drugs using a separate, independent local DP prior for each drug (Panel B.2.1), and global AE signal strengths $\lambda_{i}^{g}$ for shared by all $J=3$ drugs based on a common global DP prior (Column B.2.2). Probabilistic two-component mixtures of $\lambda_{ij}^{l}$ and $\lambda_{i}$ (Panel B.3) based on the local distribution indicators $z_{ij}^{l}$ produce the final realized signal strengths $\lambda_{ij}$ (Panel B.4).
@@ 4.2 The local-global DP Poisson model
$\lambda_{ij}=z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g};\ \lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D;\ z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
@@
Here, the realized signal strengths $\{\lambda_{1j},\dots,\lambda_{Ij}\}$ for a drug $j$ is generated by combining local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$ drawn from a drug-$j$-specific local distribution $D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$ as in (1) and global strengths $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ generated from a global distribution $D\sim\text{DP}(\alpha,G(\beta))$ shared by all drugs. The binary local distribution indicators $z_{ij}^{l}$ governed by the common local proportion $\pi^{l}$, determine the probabilistic mixing of $\lambda_{ij}^{l}$ and $\lambda_{i}^{g}$.
We note that the proposed local-global mixture DP prior (3) enables information sharing in signal strengths across both AEs and drugs. Specifically, for each drug $j$, the realized signal strengths $\{\lambda_{i^{\prime}j}:\text{all }i^{\prime}\text{ with }z_{i^{\prime}j}=1\}$ exhibit the same $D_{j}$-induced clustering (see Section 4.1) as the local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$, facilitating information sharing across AEs. Meanwhile, for any AE $i$, the signal strengths $\{\lambda_{ij^{\prime}}:\text{all }j^{\prime}\text{ with }z_{ij^{\prime}}=0\}$ form a cluster attaining the common global value $\lambda_{i}^{g}$, and enable information sharing across drugs (the global values $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ themselves admitting clustering via $D$). Figure 2B illustrates this mechanism using $I=6$ AEs and $J=3$ drugs. This bi-level clustering contrasts with the single-level (AE-only) clustering induced by the DP prior of Hu et al.hu2015signal . As demonstrated in extensive simulations (Section 5), the bi-level clustering and information sharing play a critical role in stabilizing the posterior and enhancing signal detection for $\lambda_{ij}$, particularly for drug-AE pairs where $n_{i\bullet}$ is small (rare AEs) and/or $n_{\bullet j}$ is small (newer drugs), leading to small $E_{ij}$.
@@
The Hu et al.hu2015signal model can be viewed as a “local-only” special case of our local-global model, with only drug-specific local distributions $D_{j}$ for $\lambda_{ij}$ and no global distribution (i.e., $\pi^{l}=1$). At the other extreme, a “global-only” model ($\pi^{l}=0$) yields common $\lambda_{i}$ across drugs. This is a restrictive model that cannot differentiate between drugs and only permits “global” signal detection—assessing whether an AE-$i$ is a signal for any drug under consideration (see Supplement S.1.2 for further discussion on this model).
@@
Extensive literature exists on local-global DP models for sharing information across groups while preserving DP-based clustering within each group. A common approach employs hierarchical DPs teh2004sharing , where local DPs share a global DP as a base distribution. In contrast, our approach leverages a two-component DP mixture framework antoniak1974mixtures , separately modeling local and global AE signal strengths and thus offering a more transparent bi-level clustering over AEs and drugs, enabling more flexible estimation of the local and global parameters for greater flexibility in SRS applications.
@@
In this paper, we focus on drug-based pharmacovigilance, identifying AE (row) signals given one or more drugs (columns in SRS contingency tables). An alternative AE-based approach can be performed to discover drugs strongly associated with a given set of AEs ding2020evaluation ; liu2024pattern . Our proposed approach can be readily adapted for such analyses: a row-based version of our column-based prior (3), obtained by switching the roles of $i$ and $j$, can be used for this purpose.

### v3

@@ 3.2 The Gamma Poisson Shrinker (GPS) model
where $\kappa_{ij}=\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})/[\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})+(1-\kappa)f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{2},\beta_{2})]$. Under an empirical Bayes approach, the point estimates $\hat{\kappa}$, $\hat{\alpha}_{1}$, $\hat{\beta}_{1}$, $\hat{\alpha}_{2}$, and $\hat{\beta}_{2}$ for the hyperparameters are obtained by maximizing their marginal likelihoods given $n_{ij}$ and $E_{ij}$. The resulting posterior of $\lambda_{ij}$ conditional on these estimated hyperparameters is used for inference and signal detection. Specifically, DuMouchel and Pregibon (2001) define the empirical Bayesian score $E(\log_{2}(\lambda_{ij})\mid n_{ij},E_{ij},\hat{\kappa},\hat{\alpha}_{1},\hat{\beta}_{1},\hat{\alpha}_{2},\hat{\beta}_{2})$ as a point estimate of signal strength for the drug-$j$ and AE-$i$ pair, and also suggest determining the pair to be a signal if the 5th percentile (EB05) of the above conditional posterior for $\lambda_{ij}$ exceeds a specific threshold, commonly 2. Similar to the BCPNN method, the deterministic threshold of $2$ may not necessarily ensure strong FDR control; however, a data-driven threshold controlling FDR to a pre-specified level (e.g., 0.05) may be obtained.Ahmed et al. (2010)
3.3 DP Poisson method of Hu et al.Hu et al. (2015)
Several extensions and generalizations of the two-component mixture prior used in the GPS method have been suggested in the literatureHu et al. (2015); Zhang et al. (2018); Tan et al. (2025). Of these, we focus on one of the most flexible generalizations as proposed in Hu et al.Hu et al. (2015), which employs a Dirichlet process mixture as the prior for $\lambda_{ij}$. Hu et al.Hu et al. (2015) consider the same Poisson likelihood $n_{ij}\sim$ Poisson($\lambda_{ij}E_{ij}$) as in GPS, but replace the two-gamma mixture prior for $\lambda_{ij}$ by a DP (infinite) mixture prior:
@@
Meanwhile, the atoms $\theta_{hj}$ are generated independently from the DP base gamma distribution: $\theta_{hj}\sim G(\beta_{j})$; $h\geq 1$. In practice, the infinite mixture (1) defining $D_{j}$ can be well-approximated Ishwaran and Zarepour (2002) by a finite mixture $\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$ with a sufficiently large number of components $K$; this finite mixture approximation provides a computationally tractable approach to implementing this model.
Hu et al.Hu et al. (2015) consider full Bayesian inference on $\lambda_{ij}$ by first eliciting independent vague prior distributions on the hyperparameters $\alpha_{j}$ (specifically a Uniform(0.2, 10) prior) and $\beta_{j}$ (a Uniform(0, 1) prior) and then utilizing MCMC sampling from the resulting joint posterior. Subsequently, the drug-$j$ and AE-$i$ pair is identified as a signal if the 5th posterior percentile of $\lambda_{ij}$ computed using its posterior MCMC draws exceeds a certain threshold. The authors suggest using the deterministic threshold of $2$ for this purpose; a more data-driven threshold controlling FDR at a prespecified level (e.g., $0.05$) may be obtained.Ahmed et al. (2010)
@@ 3.4 ZIP model based Pseudo LRT (frequentist) method
In parallel to Bayesian developments, major advances on SRS data mining for AE signal detection have also been made using frequentist methods, notably using the machinery of the likelihood ratio test Huang et al. (2011, 2013, 2014). These approaches often have the advantage of built-in FDR control while ensuring high sensitivity. We consider the ZIP model for the reporting count $n_{ij}\sim\text{ZIP}(n=n_{ij}\mid\lambda_{ij}\times E_{ij},\omega_{j})$. The pseudo LRT, as proposed by Chakraborty et al.Chakraborty et al. (2022), is a recent method designed to accommodate the zero-inflated nature of SRS data while easing computational demands. Similar to the standard LRT method, this approach first tests a global alternative hypothesis $H_{aj}:\lambda_{ij}>1$ for at least one $i\in\{1,\dots,I\}$ against the null hypothesis $H_{0j}:\lambda_{ij}=1$ for all $i\in\{1,\dots,I\}$ for a given drug-$j$. If $H_{aj}$ is rejected, a step-down procedure is used to conduct individual tests on $H_{aij}:\lambda_{ij}>1$.
@@
The maximum likelihood ratio statistic for drug-$j$ is then defined as $\text{MLR}_{j}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\}\end{subarray}}\text{LR}_{ij}$. In the global test, a parametric bootstrap is used to approximate the p-value $P_{j}=\Pr(\text{MLR}_{j}\geq\text{MLR}_{j}^{\text{obs}})$ using (parametric bootstrap) draws for $\text{LR}_{ij}$ under $H_{0j}$ and hence $\text{MLR}_{j}$, where $\text{MLR}_{j}^{\text{obs}}$ is the observed test statistic. If the global test is rejected, step-down p-values Chakraborty et al. (2022); Huang et al. (2011) are computed for the individual tests using the same parametric bootstrap samples for $\text{MLR}_{j}$ and using the individual observed likelihood ratio test statistic values, such that $P_{ij}=\Pr(\text{LR}_{ij}\geq\text{MLR}_{ij}^{\text{obs}})$. For multiple drugs, the same approach can be applied, defining an overall maximum likelihood ratio as $\text{MLR}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\},j\in\{1,\dots,J\}\end{subarray}}\text{LR}_{ij}$. This step-down procedure employing the same global test statistic strictly controls the false discovery rate at the same level $\alpha$ of the global test performed using the MLR Huang et al. (2011); Chakraborty et al. (2022). When multiple drugs $j$ are tested simultaneously, one considers the extended version with global maximum likelihood ratio $\max_{j\in\{1,\dots,J\}}\text{MLR}_{j}$.
@@ 4 A NOVEL LOCAL-GLOBAL DP MODEL FOR SRS DATA MINING
We begin this section by revisiting the DP Poisson model of Hu et al. Hu et al. (2015) (Section 4.1), highlighting the role of its hyperparameters in SRS signal discovery with the aid of a schematic plot (Figure 2a) and providing guidance on prior specification for these hyperparameters. We then introduce our novel local-global DP framework (Section 4.2), which accommodates between-drug associations and enables effective information sharing across drugs, thereby enhancing SRS data mining.
@@
The stick-breaking representation (2) elucidates that when $\alpha_{j}<1$, each $v_{hj}$, on average, is greater than $1/2$ with smaller values of $\alpha_{j}$ favoring $v_{hj}$ closer to one. Consequently, $\eta_{hj}$ decreases sharply in $h$, leading the mixture (1), or its finite approximation with $K$ components, to retain only a few non-empty components with sufficiently large and positive weights $\eta_{hj}$. Under a finite mixture approximation with $K_{+}<K$ non-empty components, (1) models $\lambda_{1j},\dots,\lambda_{Ij}$ using a categorical distribution with $K_{+}$ categories, placing point masses at the atoms $\theta_{1j},\dots,\theta_{K_{+}j}$ with weights $\eta_{1j},\dots,\eta_{K_{+}j}$. This fosters clustering, restricting $\lambda_{1j},\dots,\lambda_{Ij}$ to at most $K_{+}$ distinct values. When $K_{+}<I$, multiple $\lambda_{ij}$ share the same cluster $h$ admitting the common value $\theta_{hj}$, facilitating their pooling for information sharing. It follows that a smaller $\alpha$ encourages a smaller $K_{+}$, resulting in fewer clusters and thus greater information sharing among the $\lambda_{ij}$, while a larger $\alpha$ induces a larger $K_{+}$, resulting in more clusters that collectively capture more hetergeneity among the $\lambda_{i}j$ values. Importantly, when supported by the data and induced through a (moderately) large $\alpha$, values of $K_{+}>2$ are permitted in the posterior of this method. This contrasts with the GPS method of DuMouchel (1999) (see Section 3 for a review) which strictly categorizes $\lambda_{ij}$ into two groups, intuitively, signal ($\lambda_{ij}\leq 1$) and non-signal ($\lambda_{ij}\gg 1$). The DP framework, by allowing $K_{+}>2$, enhances heterogeneity acknowledgment within both signal and non-signal groups. Figure 2a illustrates this with $I=6$ AEs and $K_{+}=3$ clusters.
@@
Because a smaller $\alpha_{j}$ promotes greater information sharing, a prior assigning sufficient mass to small positive values is recommended for stable estimation of $\lambda_{ij}$. Hu et al.Hu et al. (2015) used a Uniform(0.2, 10) prior for $\alpha_{j}$. We suggest a more aggressive weakly informative default prior, viz., Exponential(rate = $\psi_{\alpha}$) with a moderately large $\psi_{\alpha}>1$. This prior favors small $\alpha_{j}$ while allowing larger values if needed and ensures conjugacy for efficient Gibbs sampling updates in the MCMC implementation. As a default choice, we suggest $\psi_{\alpha}=3$, which places $\approx 95\%$ prior mass in $(0,1)$. However, our simulation-based sensitivity analyses using typically sized SRS datasets indicate that model performance is largely robust to the choice of $\psi_{\alpha}$ when specified within a weakly informative range; see Remark Remark.
@@
Figure 2: Schematic diagrams visualizing the mechanisms of a “local only” DP prior with one drug (Panel A) and local-global mixture DP prior with $J=3$ drugs (Panel B). Throughout, darker shades represent larger values. Panel A: The observed relative reporting rates $n_{ij}/E_{ij}$ for $I=6$ AEs for the drug $j$ (Panel A.1) show high variability across $i$ (different shades of green) due to randomness in observed data. The atoms $\theta_{hj}$ (Panel A.2) of the DP prior enable clustering of $I=6$ AE signal strengths into $K_{+}=3$ non-empty DP clusters. This ensures information sharing: the AEs $(i=1,i=3,i=5)$ in cluster 1 and separately the AEs $(i=2,i=6)$ in cluster 2 (cluster memberships are visualized via arrows between Panels A.1 and A.2) share their observed $(n_{ij},E_{ij})$ information to inform their cluster-specific DP atoms (single atom per cluster). These DP atoms, together with the cluster memberships, produce the final $I=6$ AE signal strengths $\lambda_{ij}$ (Panel A.3). Panel B: The noisy observed data $(n_{ij},E_{ij})$ leading to noisy observed relative reporting rates $n_{ij}/E_{ij}$ (Panel B.1; different shades of green) are used to produce “local” signal strengths $\lambda_{ij}^{l}$ for $I=6$ AEs and $J=3$ drugs using a separate, independent local DP prior for each drug (Panel B.2.1), and global AE signal strengths $\lambda_{i}^{g}$ for shared by all $J=3$ drugs based on a common global DP prior (Column B.2.2). Probabilistic two-component mixtures of $\lambda_{ij}^{l}$ and $\lambda_{i}$ (Panel B.3) based on the local distribution indicators $z_{ij}$ produce the final realized signal strengths $\lambda_{ij}$ (Panel B.4).
@@ 4.2 The local-global DP Poisson model
$\lambda_{ij}=z_{ij}\lambda_{ij}^{l}+(1-z_{ij})\lambda_{i}^{g};\ \lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D;\ z_{ij}\sim\text{Bernoulli}(\pi)$
@@
Here, the realized signal strengths $\{\lambda_{1j},\dots,\lambda_{Ij}\}$ for a drug $j$ is generated by combining local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$ drawn from a drug-$j$-specific local distribution $D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$ as in (1) and global strengths $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ generated from a global distribution $D\sim\text{DP}(\alpha,G(\beta))$ shared by all drugs. The binary local distribution indicators $z_{ij}$ governed by the common local proportion $\pi$, determine the probabilistic mixing of $\lambda_{ij}^{l}$ and $\lambda_{i}^{g}$.
We note that the proposed local-global mixture DP prior (3) enables information sharing in signal strengths across both AEs and drugs. Specifically, for each drug $j$, the realized signal strengths $\{\lambda_{i^{\prime}j}:\text{all }i^{\prime}\text{ with }z_{i^{\prime}j}=1\}$ exhibit the same $D_{j}$-induced clustering (see Section 4.1) as the local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$, facilitating information sharing across AEs. Meanwhile, for any AE $i$, the signal strengths $\{\lambda_{ij^{\prime}}:\text{all }j^{\prime}\text{ with }z_{ij^{\prime}}=0\}$ form a cluster attaining the common global value $\lambda_{i}^{g}$, and enable information sharing across drugs (the global values $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ themselves admitting clustering via $D$). Figure 2B illustrates this mechanism using $I=6$ AEs and $J=3$ drugs. This bi-level clustering contrasts with the single-level (AE-only) clustering induced by the DP prior of Hu et al.Hu et al. (2015). As demonstrated in extensive simulations (Section 5), the bi-level clustering and information sharing play a critical role in stabilizing the posterior and enhancing signal detection for $\lambda_{ij}$, particularly for drug-AE pairs where $n_{i\bullet}$ is small (rare AEs) and/or $n_{\bullet j}$ is small (newer drugs), leading to small $E_{ij}$.
@@
As a special case of our local-global model, the “local-only” DP model of Hu et al.Hu et al. (2015) can be recovered by setting the common local proportion $\pi=1$, which leads to $z_{ij}=1$ and hence $\lambda_{ij}=\lambda_{ij}^{l}$ for all $i=1,\dots,I$ and $j=1,\dots,J-1$. Conversely, a “global-only” model can be obtained by setting $\pi=0$, leading to $\lambda_{ij}=\lambda_{i}^{g}$ for all $i$ and all $j<J$. We note, however, that this model is considerably more restrictive, as it does not distinguish among $\{\lambda_{i1},\dots,\lambda_{i,J-1}\}$ for any $i$. This yields the simplified likelihood $n_{ij}\sim\text{Poisson}(\lambda_{i}E_{ij})$, which produces the sufficient statistic $\sum_{j=1}^{J-1}n_{ij}\sim\text{Poisson}(\lambda_{i}\sum_{j=1}^{J-1}E_{ij})$ for $\lambda_{i}$. The model reduction also impacts signal detection. Under both the general local-global and local-only frameworks, hypothesis testing is conducted for each individual $\lambda_{ij}$. In contrast, under the global-only model, inference reduces to testing whether $\lambda_{i}^{g}>1$. Testing $\lambda_{i}^{g}>1$ amounts to assessing whether the aggregated count $\sum_{j=1}^{J-1}n_{ij}$ is large relative to the expected total count $\sum_{j=1}^{J-1}E_{ij}$ under the reduced model. This represents a type of “global test” that collapses all drugs into one homogeneous group, whereas the local-global (and local-only) models allow hypothesis testing for each individual drug-AE pair. Consequently, direct comparison between the global-only and local-global models is usually not practically meaningful, as they target different inferential objectives.
@@
Extensive literature exists on local-global DP models for sharing information across groups while preserving DP-based clustering within each group. A common approach employs hierarchical DPs Teh et al. (2004), where local DPs share a global DP as a base distribution. In contrast, our approach leverages a two-component DP mixture framework Antoniak (1974), separately modeling local and global AE signal strengths and thus offering a more transparent bi-level clustering over AEs and drugs, enabling more flexible estimation of the local and global parameters for greater flexibility in SRS applications.
@@
In this paper, we focus on drug-based pharmacovigilance, identifying AE (row) signals given one or more drugs (columns in SRS contingency tables). An alternative AE-based approach can be performed to discover drugs strongly associated with a given set of AEs Ding et al. (2020); Liu et al. (2024). Our proposed approach can be readily adapted for such analyses: a row-based version of our column-based prior (3), obtained by switching the roles of $i$ and $j$, can be used for this purpose.